OpenStudy (lovelyharmonics):

more continuity c:

OpenStudy (lovelyharmonics):

Find the indicated limit, if it exists.

OpenStudy (anonymous):

So, whats the limit as we approach 0 from the left? Which part of the piece-wise function would we use?

OpenStudy (lovelyharmonics):

i have no idea. i cant even figure out how to put this into my calculator .-.

OpenStudy (anonymous):

So, if we approach zero from the left, we're approaching when x is negative. If x is negative, which part of the function would we use?

OpenStudy (lovelyharmonics):

the second part

OpenStudy (anonymous):

No, the first. That notation means: When x "is less than" 0, f(x) = 5x-8 When x "is greater than or equal to" 0, f(x) = |-4-x| So, when approaching from the left (from the negative side), we'll use the first equation. Since there isn't anything that would stop us from having x=0 in that equation (we aren't dividing by x anywhere, for example), the limit is just what you get when you plug in 0 for x. \[\lim_{x \rightarrow 0^-}f(x) = 5(0)-8 = -8\] So, the left-hand limit is -8. What would we do to find the limit from the right hand side?

OpenStudy (lovelyharmonics):

itd be |-4-(0)| =|-4|=4

OpenStudy (anonymous):

Yep! Are the two limits the same? If they are, then the limit exists, and that's the answer. If they are not, then the limit does not exist.

OpenStudy (lovelyharmonics):

they dont match so they arnt the same right?

OpenStudy (anonymous):

Correct. So the limit does not exist.

OpenStudy (lovelyharmonics):

okay thank you c:

OpenStudy (anonymous):

Does it make sense?